# Turning to Online Forums for Legal Information: A Case Study of GDPR’s Legitimate Interests

### v1

@@ Abstract
Practitioners building online services and tools often turn to online forums such as Reddit, Law Stack Exchange, and Stack Overflow for legal guidance to ensure compliance with the GDPR. The legal information presented in these forums directly impact present-day industry practitioner’s decisions. Online forums can serve as gateways that, depending on the accuracy and quality of the answers provided, may either support or undermine the protection of privacy and data protection fundamental rights. However, there is a need for deeper investigation into practitioners’ decision-making processes and their understanding of legal compliance.
@@
Our analysis found that applying the “legitimate interests” legal basis is complex for practitioners, with important implications for how the GDPR is implemented in practice. The legal analysis showed that crowdsourced legal information tends to be legally sound, though sometimes incomplete. We outline recommendations to improve the quality of online forums by ensuring that responses are more legally sound and comprehensive, enabling practitioners to apply legitimate interests effectively in practice and uphold the GDPR.
@@ 1 Introduction
The General Data Protection Regulation (GDPR) came into effect in May 2018 to protect the personal data of users residing within the European Union [28]. However, it also brought challenges for organizations who must translate abstract legal principles into specific technical requirements to assure compliance [18, 36, 42, 44, 70].
@@
How then do practitioners (anyone involved in applying the GDPR for their respective organization, including developers, designers, business owners, and others), particularly those without in-house legal support, navigate GDPR obligations? While some official compliance guidance by regulatory authorities exists, some practitioners, also referred to as posters in this paper [^1] Here, posters refer to the practitioners who are posting questions in online forums. We also use the term practitioners when referring to the same users posting questions too., may turn to online forums to seek legal information. We use the term “legal information” [^2] There is a difference between “legal information” and “legal advice.” When posters refer to their concrete cases on Stack Overflow, they are provided with legal information tuned to their context. Legal advice is instead given by a legal professional and establishes an attorney-client relationship, which is not the case in online forums. In this paper, we use the term “information” to signify the legal case-related recommendations about a given post. to signify the legal case-related recommendations in a given post. Online forums can provide scenario-specific, cost-free, and quick access to answers, and their use by industry practitioners has been documented in other contexts [37, 74, 81]. Online forums can serve as gateways that, depending on the accuracy and quality of the answers provided, may either support or undermine the protection of privacy and data rights through adherence to data protection regulations.
In this paper, we investigate online forums in the context of GDPR compliance. We look at how practitioners seek legal information using online forums (Reddit, Law Stack Exchange, and Stack Overflow) to be compliant with the GDPR. Not only does the legal information presented in these forums directly impact present-day industry practitioner’s decisions, in the near future, it is possible that more practitioners and users may start turning to LLMs and other AI tools for legal information [6] which are trained on online forum data [41]. Incorrect information present in these forums will thus have even bigger consequences with the usage of these new technologies [13]. Therefore, it is important that online forums provide correct legal information for current and potential future uses of legal information.
@@
The GDPR offers an extensive data protection regime that yields thousands of posts and responses when searched for in online forums. To make data analysis more feasible, we decided to focus our investigation specifically on the questions and responses that practitioners posed relating to “legitimate interests,” one of the GDPR’s six legal grounds for collecting data. We focused on legitimate interest due to the understudied, yet highly important nature of this legal ground due to its flexibility and breadth, covering legal bases for data collection not currently mentioned by the GDPR [35].
@@
Legitimate interest is defined in Art. 6(1)(f) of the GDPR as the processing that is necessary for the legitimate interests pursued by data controllers [68]) or third parties [28, 30]. While consent is generally well-established and recognized due to the proliferation of cookie banners, legitimate interest is not [48]. In this paper, we analyze online forum discussions relating to legitimate interests to understand what confusions practitioners may have in applying it, and assumptions practitioners have about this legal basis.
@@
Contributions. Our paper presents two main contributions. First, through a qualitative analysis of the questions in the online forums, we identified the main challenges practitioners may face when applying the legitimate interest legal basis ‘in the wild” by studying online forum posts. Second, our analysis of forum answers offering legal information indicates that while information may seem legally sound, there are precautions practitioners must take before relying on these forums. In our paper, we offer recommendations for how online legal forum posts could be formulated to ensure legally sound answers, therefore helping to better protect users’ fundamental rights to privacy and data protection.
@@ 2.2 Importance of Online Forums for Software Development
Online forums are an important source of information for software developers [9, 11, 60]. Previous research has shown that Stack Overflow influences developers’ practices, such as code reuse from Stack Overflow posts [81], and impacts developers’ productivity on GitHub [77]. However, they may also promote poor practices such as reusing code, which has negative security and privacy impacts [37].
The number of GDPR-related Stack Overflow posts in recent years has been increasing [75], which suggests that privacy ad data protection are becoming more important for developers [72]. The knowledge shared in online forums can impact how developers abide to security and data protection obligations [4, 37, 50, 75]. Privacy-related discussions on Stack Overflow seem to be connected to external events relevant to privacy, such as new privacy restrictions [72, 50]. When such an event happens, developers often felt privacy restrictions required more efforts from them without many benefits [50].
@@ 2.2.1 Online legal compliance information
Several online resources have emerged, aimed at helping industry practitioners, often developers, apply the GDPR. Some are from Data Protection Authorities (DPAs), such as France’s DPA “GDPR Developer’s Guide” [^3] https://www.cnil.fr/en/gdpr-developers-guide.. Other sources of online information derive from developers aimed at helping fellow developers. [^4] https://gdpr4devs.com/ There are also GDPR compliance guides from Consent Management Platforms (CMPs), such as OneTrust. [^5] https://www.onetrust.com/blog/gdpr-compliance/
@@ 3 Methods
As the topic of our analysis spans questions at the intersection of technology and law, we collected data from three popular online forums that technology practitioners may turn to for advice: Reddit, Law Stack Exchange, and Stack Overflow. We collected and analyzed this data to understand the struggles and discussions practitioners are having online when it comes to GDPR compliance, and to evaluate the legal soundness of the answers.
@@
We searched for posts and comments from industry practitioners containing the phrase “legitimate interest(s)” on September 27, 2024. We made requests to the Reddit API to collect relevant data for our analysis, filtering results based on the subreddit, and stored this data in a CSV file. Our scrape collected the title of the post, URL (in case further analysis was needed), comments, the post’s body text, subreddit name, upvotes for the post, age of the post, and word count. We manually removed some posts from our dataset which were out of the scope of this study (e.g., describing legitimate interests in offline contexts). Most of our posts came from the r/GDPR and r/privacy subreddit due to the relevance of these posts for our analysis.
@@ 3.1 Dataset
Ethical Considerations of Using Online Forum Data. A study using social media data was preferred over other research methods, such as interviewing or surveying industry practitioners, because: i) it would be difficult to gather industry participants who would be willing to discuss their company’s data practices in a research context, ii) this data is readily available for analysis, and iii) online forums allow researchers to view how practitioners might navigate legal compliance “in the wild”. A social media analysis would ensure a more targeted and scalable study of what industry practitioners are discussing for technical implementations of legitimate interest. Reddit and Stack Exchange (including Stack Overflow) are two commonly-used sources of data for academic research [65, 34].
@@ 4 Analysis 1: Practitioner Discussions and Inquiries
Qualitative Analysis. We analyzed the content from the three forums and conducted a thematic analysis to find underlying patterns within the dataset by labeling important parts in the data (codes), and grouping these codes to form themes [17]. To investigate what practitioners discuss in online forums, we had two annotators analyse the collected posts. They first annotated a random set of 15 posts together to discuss what to focus the analysis upon. They decided to keep track of: i) the (perceived) role or industry of the poster, ii) the general topic they were seeking information about, and iii) codes related to the poster’s query. After this was established, the annotators coded the same random set of 15 posts to calculate the interrater reliability. The agreement rate was 85%, with a Cohen’s $\kappa=$ 0.66, which indicates substantial agreement [54]. Due to the high agreement, the annotators split the dataset, and annotated the remaining half (107 posts) on their own. In total, 135 codes were identified, which fit into nine major themes. When describing our results only paraphrased quotes are included, sometimes with ethically fabricated information to maintain anonymity [52].
@@ 4.1 Results: Practitioner Discussions and Inquiries
Applying legitimate interests, and more broadly, the GDPR, is challenging for various types of practitioners. Often, posters explicitly indicated their role. However, other times we were able to infer their role based on the nature of the question. For example, we assumed that technical questions were posed by a developer, and client communication questions came from someone working within the marketing domain. The majority of online GDPR compliance information tends to target developers, but our analysis showed that other roles are involved, or have an interest, in the application of legitimate interests.
@@
Deciding on the roles responsible for GDPR compliance is important for the concrete application of legitimate interests, since the lawful application of this legal basis has been doubted in previous research [48, 35], therefore impacting the GDPR compliance for that company. We found that practitioners often queried about who in their company is responsible for GDPR compliance. Oftentimes, posters assumed it was the role of the developer to be the “de facto” data protection officer for their company, and some others believed it was the role of the owner of a given service. The following quote reveals why developers are often given the responsibility: “I’m a developer, not a lawyer, who is trying to comply with the GDPR and ePD as best as I can for my company. If we had more money, I would hire a lawyer to handle this for me, but we don’t so I’m in charge of this.” Discerning this role clearly is crucial since compliance obligations and liability depend on their accurate characterization (Recital 79 GDPR) [43, 29]. Given the broad and flexible nature of legitimate interest legal basis [45, 35], it is especially important to decide who is accountable for GDPR compliance within a team for ensuring the proper application of this legal basis.
@@
Fourth, many posts held the assumption that legitimate interest may be applied to retain data in a scenario where users exercise the right to be forgotten (Art. 17 GDPR). This has practical implications because controllers can lose money if data is deleted, and some companies may face challenges locating the user’s data upon an account deletion request, according to our analysis.
@@ 4.1.4 Questions from practitioners indicate knowledge, but also misunderstandings of the GDPR
Trying to comply with the GDPR is made more difficult when it comes to the legitimate interest legal ground, as these are more broad and flexible in their application, with little guidance on how they should be used. While other legal bases such as consent are well-established with the advent of cookie banners and the general understanding that companies need to request user consent in order to collect and process data, there is less knowledge about the legitimate interest legal ground [48]. On the other hand, we find that developers (and those applying the GDPR) were often confused by their own service’s data processing. The examples below illustrate commons points of confusion.
@@
Previous work has showed that the legitimate interest ground is commonly seen in cookie banners, being treated similarly to consent [48, 53]. Therefore, the challenges practitioners face in distinguishing whether users have provided valid consent indicate that companies may struggle to manage and categorize the data collected from users and align it with the appropriate legal basis.
Anonymization and pseudonymization are misunderstood. Oftentimes, it was mentioned that true anonymization is difficult, and just because data is hashed, it does not mean it can evade GDPR rquirements. However, there was a lot of discussion about whether anonymization and pseudonymization would be sufficient to protect user’s personal data, and whether the GDPR applies to anonymized and pseudonymized data: “I believe the GDPR wouldn’t apply if I properly anonymize the data because then it wouldn’t count as personal information.”
@@ 5.1.1 Completeness
However, in our sample, no regulatory decisions from DPAs were cited (which include concrete cases that could inform several questions and answers closely related to the posts), even if there exists a publicly available online repository with relevant DPA decisions and more general GDPR insights that can be shared across Europe. [^6] https://gdprhub.eu/
@@ 5.1.2 Soundness
While the legal basis of legitimate interest is open-ended (with a broad and unspecific scope), several regulatory decisions and EDPB guidelines determined that personalized advertising purposes cannot rely on legitimate interest grounds [1, 5, 30, 53, 25]. Moreover, some legal scholars argued that legitimate interests should not be used as a ground to collect data by higher education institutions for proctoring [38]. This reveals that there is a mismatch between the the legal field and some commentators’ conceptions of legitimate interests, which may result in this legal ground being misused if posters were to follow this legal information.
Legitimate interests are incorrectly invoked under the Right to be Forgotten. Several posters wanted to know if they could deny a request for account deletion from their users under legitimate interests. They moreover wondered whether they could keep data based on legitimate interests in cases where relevant data records could not be located. Answers related to these posts mentioned the right to be forgotten, but did not account for all the legal grounds for this right to be exercised (under Article 17(1)), or its exceptions (prescribed in Article 17(3) and Recital 65) that could be potentially applicable to override the deletion request. We argue that incomplete answers about erasure obligations are mostly due to the difficulty of operationalizing such grounds and exceptions in practice [8].
@@
Legitimate interests are incorrectly invoked in the context of publicly available data. In response to posts asking about whether they can use publicly available data for their commercial purposes, some commentators asserted that scraping data publicly available on the web could be based on legitimate interests and it would suffice to disclose the practice in a privacy policy on a relevant website, as per Article 14 GDPR.
However, the reuse of publicly available online data for commercial purposes requires user consent [20, 19, 22]. The answers also did not account for potential sensitive categories of data, nor the obligation to directly and actively inform the involved data subjects about their rights (e.g., to object to legitimate interests), the source of the scraped data, as well as the mandatory information required by Article 14(1-3) GDPR [32, 23]. Yet, some answers acknowledged that the public nature of the scraped data might be factored in the legitimate interest test: “The data subject making this info public is a factor in the legitimate interest balancing test, but it depends on how you use the data and if they can expect their data to be used this way.”

### v2

@@ Abstract
Practitioners building online services and tools often turn to online forums such as Reddit, Law Stack Exchange, and Stack Overflow for legal guidance to ensure compliance with the GDPR. The legal information presented in these forums directly impacts present-day industry practitioner’s decisions. Online forums can serve as gateways that, depending on the accuracy and quality of the answers provided, may either support or undermine the protection of privacy and data protection fundamental rights. However, there is a need for deeper investigation into practitioners’ decision-making processes and their understanding of legal compliance when seeking for legal information online.
@@
Our analysis found that applying the legal basis of legitimate interest is complex for practitioners, with important implications for how the GDPR is implemented in practice. The legal analysis showed that crowdsourced legal information tends to be legally sound, though sometimes incomplete. We outline recommendations to improve the quality of online forums by ensuring that responses are more legally sound and comprehensive, enabling practitioners to apply legitimate interests effectively in practice and uphold the GDPR.
@@ 1 Introduction
The General Data Protection Regulation (GDPR) came into effect in May 2018 to protect EU users’ personal data [32]. However, it also brought challenges for organizations who must translate abstract legal principles into specific technical requirements to assure compliance [20, 42, 48, 51, 79].
@@
How then, do practitioners (anyone involved in applying the GDPR for their respective organization, including developers, designers, business owners, and others), particularly those without in-house legal support, navigate GDPR obligations? While some official compliance guidance by regulatory authorities exists, some practitioners, also referred to as posters in this paper [^1] Here, posters refer to the practitioners who are posting questions in online forums. We also use the term practitioners when referring to the same users posting questions too., may turn to online forums to seek legal information. We use the term “legal information” [^2] There is a difference between “legal information” and “legal advice.” When posters refer to their concrete cases on Stack Overflow, they are provided with legal information tuned to their context. Legal advice is instead given by a legal professional and establishes an attorney-client relationship, which is not the case in online forums. In this paper, we use the term “information” to signify the legal case-related recommendations about a given post. to signify the legal case-related recommendations in a given post. Online forums can provide scenario-specific, cost-free, and quick access to answers, and their use by industry practitioners has been documented in other contexts [43, 84, 92]. Online forums can serve as gateways that, depending on the accuracy and quality of the answers provided, may either support or undermine the protection of privacy and data rights through adherence to data protection regulations.
In this paper, we investigate online forums in the context of GDPR compliance. We look at how practitioners seek legal information using the following online forums – Reddit, Law Stack Exchange, and Stack Overflow – to be compliant with the GDPR.
We explore online forums for the following reasons. First, developers tend to primarily use online forums to support their privacy-related decisions [69, 85]. Developers often struggle to meet legal requirements as they often lack privacy and legal expertise [48, 49]. Second, other tools to support their decision-making processes, such as official and reliable guidance for developers, are largely lacking [23, 36]. These guides are not contextually-sensitive for developer’s projects, thus they are left with industry-based sources [86] which, while might be practical, are typically unofficial, potentially biased, and not subject to rigorous legal or regulatory review. Consequently, the legal information presented in these forums directly impacts present-day industry practitioner’s decisions, given the prevalence of forums to aid developers [85]. With the development of Large-Language Models (LLMs), a novel source to access online legal information has been introduced into the software development process. This shift is reflected in trends, such as Stack Overflow reporting a decline in user activity, which they attributed to the rise of AI systems [25]. Third, more practitioners and users have started using LLMs and other AI assistants for legal information seeking and use [7, 25] which are trained on online forum data [47]. Incorrect information present in these forums will thus have even bigger consequences with the usage of these new technologies [15]. Therefore, it is important that online forums provide correct legal information for current and potential future uses of legal information.
@@
The GDPR offers an extensive data protection regime that yields thousands of posts and responses when searched for in online forums. To make data analysis more feasible, we decided to focus our investigation specifically on the questions and responses that practitioners posed relating to “legitimate interests,” one of the GDPR’s six legal grounds for collecting data. We focused on legitimate interest because of its vague nature, yet is highly important due to its flexibility and breadth, covering legal bases for data collection not currently mentioned by the GDPR [41].
@@
Legitimate interest is defined in Art. 6(1)(f) of the GDPR as the processing that is necessary for the legitimate interests pursued by data controllers [77] or third parties [32, 34]. While consent is generally well-established and recognized due to the proliferation of cookie banners, legitimate interest is not [55]. In this paper, we analyze online forum discussions relating to legitimate interests to understand what confusions practitioners may have in applying it, and assumptions practitioners have about this legal basis.
@@
Contributions. Our paper presents two main contributions. First, through a qualitative analysis of the questions in the online forums, we identified the main challenges practitioners may face when applying the legitimate interest legal basis “in the wild” by studying online forum posts. Second, our analysis of forum answers offering legal information indicates that while information may seem legally sound, there are precautions practitioners must take before relying on these forums. In our paper, we offer recommendations for how online legal forum posts could be formulated to ensure legally sound answers, therefore helping to better protect users’ fundamental rights to privacy and data protection.
@@ 2.2 Importance of Online Forums for Software Development
Online forums are an important source of information for software developers [10, 12, 68]. Previous research has shown that Stack Overflow influences developers’ practices, such as code reuse from Stack Overflow posts [92], and impacts developers’ productivity on GitHub [88]. However, they may also promote poor practices such as reusing code, which has negative impacts on the security and privacy of systems [43].
The number of GDPR-related Stack Overflow posts in recent years has been increasing [85], which suggests that privacy and data protection are becoming more important for developers [82]. The knowledge shared in online forums can impact how developers abide to security and data protection obligations [5, 43, 57, 85]. Privacy-related discussions on Stack Overflow seem to be connected to external events relevant to privacy, such as new privacy restrictions [82, 57]. When such an event happens, developers often felt privacy restrictions required more efforts from them without many benefits [57].
@@ 2.2.1 Online legal compliance information
Several online resources have emerged, aimed at helping industry practitioners, often developers, apply the GDPR. Some are from Data Protection Authorities (DPAs), such as the French DPA’s “GDPR Developer’s Guide” [23] and European Data Protection Board’s “Data Protection Guide for Small Business” [36]. Other sources of online information derive from developers aimed at helping fellow developers [4], and GDPR compliance guides from Consent Management Platforms (CMPs), such as OneTrust [86].
@@ 3 Methods
As the topic of our analysis spans questions at the intersection of technology and law, we collected data from three popular online forums that technology practitioners may turn to for advice: Reddit, Law Stack Exchange, and Stack Overflow. Our general approach was to collect and analyze this data to understand the struggles and discussions practitioners are having online when it comes to GDPR compliance, and to evaluate the legal soundness of the answers.
@@ 3.1 Data Collection
We searched for posts and comments from industry practitioners containing the phrase “legitimate interest(s)” on September 27, 2024. We made requests to the Reddit API to collect relevant data for our analysis, filtered results based on the subreddit, and stored this data in a CSV file. Our scrape collected the title of the post, URL (in case further analysis was needed), comments, the post’s body text, subreddit name, upvotes for the post, age of the post, and word count. We manually removed some posts from our dataset which were out of the scope of this study (e.g., describing legitimate interests in offline contexts). Most of our posts came from the r/GDPR and r/privacy subreddit due to the relevance of these posts for our analysis.
@@ 3.2 Dataset
Ethical Considerations of Using Online Forum Data. Social media data can be rich and allow researchers to understand social phenomena, but also presents several ethical challenges [13]. A study using social media data was preferred over other research methods, such as interviewing or surveying industry practitioners, because: i) it would be difficult to gather industry participants who would be willing to discuss their company’s data practices in a research context, ii) this data is readily available for analysis, and iii) online forums allow researchers to view how practitioners might navigate legal compliance “in the wild”. A social media analysis would ensure a more targeted and scalable study of what industry practitioners are discussing for technical implementations of legitimate interest. Reddit and Stack Exchange (including Stack Overflow) are two commonly-used sources of data for academic research [74, 40].
@@ 4 Analysis 1: Practitioner Discussions and Inquiries
Qualitative Analysis. We analyzed the content from the three forums and conducted a thematic analysis to find underlying patterns within the dataset by labeling important parts in the data (codes), and grouping these codes to form themes [19]. To investigate what practitioners discuss in online forums, we had two annotators analyse the collected posts. They first annotated a random set of 15 posts together to discuss what to focus the analysis upon. They decided to keep track of: i) the (perceived) role or industry of the poster, ii) the general topic they were seeking information about, and iii) codes related to the poster’s query. After this was established, the annotators coded the same random set of 15 posts to calculate the interrater reliability. The agreement rate was 85%, with a Cohen’s $\kappa=$ 0.66, which indicates substantial agreement [61]. Due to the high agreement, the annotators split the dataset, and annotated the remaining half (107 posts) on their own. In total, 135 codes were identified from the dataset, which fit into nine major themes. When describing our results only paraphrased quotes are included, sometimes with ethically fabricated information to maintain anonymity [59].
@@ 4.1 Results: Practitioner Discussions and Inquiries
Applying legitimate interests, and more broadly, the GDPR, is challenging for various types of practitioners. Often, posters explicitly indicated their role. However, other times we were able to infer their role based on the nature of the question. For example, we assumed that technical questions were posed by a developer, and client communication questions came from someone working within the marketing domain. The majority of online GDPR compliance information tends to target developers, but our analysis showed that other roles are involved, or have an interest, in the application of legitimate interests, in line with recent work [81].
@@
Deciding on the roles responsible for GDPR compliance is important for the concrete application of legitimate interests, since the lawful application of this legal basis has been doubted in previous research [55, 41], therefore impacting the GDPR compliance for that company. We found that practitioners often queried about who in their company is responsible for GDPR compliance. Oftentimes, posters assumed it was the role of the developer to be the “de facto” data protection officer for their company, and some others believed it was the role of the owner of a given service. The following quote reveals why developers are often given the responsibility: “I’m a developer, not a lawyer, who is trying to comply with the GDPR and ePD as best as I can for my company. If we had more money, I would hire a lawyer to handle this for me, but we don’t so I’m in charge of this.” Discerning this role clearly is crucial since compliance obligations and liability depend on their accurate characterization (Recital 79 GDPR) [50, 33]. Given the broad and flexible nature of legitimate interest as a legal basis [52, 41], it is especially important to decide who is accountable for GDPR compliance within a team for ensuring the proper application of this legal basis.
@@
Fourth, many posts held the assumption that legitimate interest may be applied to retain data in a scenario where users exercise the “Right to be Forgotten” (Art. 17 GDPR). This has practical implications because controllers can lose money if data is deleted, and some companies may face challenges locating the user’s data upon an account deletion request, according to our analysis.
@@ 4.1.4 Questions from practitioners indicate knowledge, but also misunderstandings of the GDPR
Trying to comply with the GDPR is made more difficult when it comes to the legitimate interest legal ground, as these are more broad and flexible in their application, with little guidance on how they should be used. While other legal bases such as consent are well-established with the advent of cookie banners and the general understanding that companies need to request user consent in order to collect and process data, there is less knowledge about the legitimate interest legal ground [55]. However, we find that developers (and those applying the GDPR) were often confused by their own service’s data processing. The examples below illustrate commons points of confusion.
@@
Previous work has shown that the legitimate interest ground is commonly seen in cookie banners, being treated similarly to consent [55, 60]. Therefore, the challenges practitioners face in distinguishing whether users have provided valid consent indicate that companies may struggle to manage and categorize the data collected from users and align it with the appropriate legal basis.
Anonymization and pseudonymization are misunderstood. Oftentimes, it was mentioned that true anonymization is difficult, and just because data is hashed, it does not mean it can evade GDPR requirements. However, there was discussion about whether anonymization and pseudonymization would be sufficient to protect user’s personal data, and whether the GDPR applies to anonymized and pseudonymized data: “I believe the GDPR wouldn’t apply if I properly anonymize the data because it wouldn’t count as personal information.”
@@ 5.1.1 Completeness
However, in our sample, no regulatory decisions from DPAs were cited (which include concrete cases that could inform several questions and answers closely related to the posts), even if there exists a publicly available online repository with relevant DPA decisions and more general GDPR insights that can be shared across Europe [63].
@@ 5.1.2 Soundness
While the legal basis of legitimate interest is open-ended (with a broad and unspecific scope), several regulatory decisions and EDPB guidelines determined that personalized advertising purposes cannot rely on legitimate interest grounds [1, 6, 34, 60, 29]. Moreover, some legal scholars argued that legitimate interests should not be used as a ground to collect data by higher education institutions for proctoring [44]. This reveals that there is a mismatch between the legal field and some commentators’ conceptions of legitimate interests, which may result in this legal ground being misused if posters were to follow this legal information.
Legitimate interests are incorrectly invoked under the “Right to be Forgotten”. Several posters wanted to know if they could deny a request for account deletion from their users under legitimate interests. They moreover wondered whether they could keep data based on legitimate interests in cases where relevant data records could not be located. Answers related to these posts mentioned the right to erasure, but did not account for all the legal grounds for this right to be exercised (under Article 17(1)), or its exceptions (prescribed in Article 17(3) and Recital 65) that could be potentially applicable to override the deletion request. We argue that incomplete answers about erasure obligations are mostly due to the difficulty of operationalizing such grounds and exceptions in practice [9].
@@
Legitimate interests are incorrectly invoked in the context of publicly available data. In response to posts asking about whether they can use publicly available data for their commercial purposes, some commentators asserted that scraping data publicly available on the web could be based on legitimate interests and it would suffice to disclose the practice in a privacy policy on a relevant website, as per Article 14 GDPR. The reuse of publicly available online data for commercial purposes requires user consent in most data processing scenarios [24, 37, 21, 26]. The legitimate interest basis can only in theory justify reuse of this data if the the three-tier test is fulfilled and most importantly, considers large-scale data collection, user’s reasonable expectations of their data being reused , and inclusion of sensitive data, requirements that ultimately impede the use of this legal grounding.
The answers also did not account for potential sensitive categories of data, nor the obligation to directly and actively inform the involved data subjects about their rights (e.g., to object to legitimate interests), the source of the scraped data, as well as the mandatory information required by Article 14(1-3) GDPR [38, 27]. Yet, some answers acknowledged that the public nature of the scraped data might be factored in the legitimate interest test: “The data subject making this info public is a factor in the legitimate interest balancing test, but it depends on how you use the data and if they can expect their data to be used this way.”